OpenStudy (lstrasmann):

Given the two triangles are similar, what is the value of x? Round to the nearest hundredth.

OpenStudy (mathstudent55):

Figure?

OpenStudy (lstrasmann):

A. x = -1.67 B. x = - 0.833 C. x = 0.833 D. x = 6.67

OpenStudy (lstrasmann):

Sorry about that, I couldn't figure out where the pic was saved.

OpenStudy (mathstudent55):

The lengths of corresponding sides of similar triangles are proportional. Side 20 corresponds to side 8. Side x + 10 corresponds to side x. The ratio 20/8 is equal to the ratio (x + 10)/x. 20/8 = (x + 10)/x Solve the above equation for x. Start by cross multiplying.

OpenStudy (lstrasmann):

\[\frac{ 20 }{ 8 }=\frac{ (x + 10) }{ x }\]\[20x = 8(x+10)\]\[20x=8x+80\]\[12x=80\]\[x \approx 6.67\]So it's just a proportion?

OpenStudy (lstrasmann):

Can you help me with two more?

OpenStudy (mathstudent55):

Correct. Yes, because of the statement above about similar triangles.

OpenStudy (lstrasmann):

ΔSWK ~ ΔEBT. Find the value of x and y. A. x = 30; y = 62 B. x = 62; y = 30 C. x = 62; y = 40 D. x = 110; y = 40